OpenStudy (anonymous):

what is the name of this ^2

OpenStudy (anonymous):

Exponent. Specifically here to the power of 2.

OpenStudy (ghazi):

exponent, power , radical

OpenStudy (anonymous):

\[5^2 = 5 *5 = 25\]